Circular No. 7281

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1999 T2 (LINEAR) Additional astrometry (including prediscovery observations from Sept.

25 and Oct. 5) appear on MPEC 1999-T64, from whence the following parabolic orbital elements are taken:

T = 2000 Nov. 24.6865 TT  ω = 104.^{o}9646
Ω = 14.8207 }2000.0
q = 3.022536 AU  i = 110.9028
1999TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
Oct. 9 23^{h}05.^{m}65 +29^{o}41.^{′}7 4.062 4.912 144.^{o}8 6.^{o}7 16.0
19 22 52.55 +29 00.7 4.060 4.844 137.8  7.9 15.9
29 22 40.93 +28 10.5 4.088 4.776 129.0  9.3 15.8
Nov. 8 22 31.15 +27 16.2 4.143 4.709 119.4 10.6 15.8
18 22 23.39 +26 23.0 4.216 4.642 109.6 11.6 15.8
28 22 17.66 +25 35.5 4.302 4.576  99.9 12.3 15.8
Dec. 8 22 13.83 +24 56.7 4.394 4.510  90.4 12.6 15.8
18 22 11.72 +24 29.0 4.486 4.444  81.2 12.6 15.7

GRB 991014 G. Tassone, BeppoSAX Science Operations Center, Telespazio, Rome;

J. in ’t Zand, Space Research Organization Netherlands, Utrecht; F. Frontera, Istituto Tecnologie e Studio Radiazioni Extraterrestri, Bologna; and G. Gandolfi, Istituto di Astrofisica Spaziale, CNR, Rome, report: “A relatively short γ-ray burst triggered the BeppoSAX Gamma-Ray Burst Monitor (GRBM) on Oct. 14.91148 UT and was detected in the Wide Field Camera unit 2. The burst lasted for ∼ 3 s at 40–700 keV and 10 s at 2–26 keV. The peak flux is 2.5 Crab at 2–26 keV, and the peak photon countrate as measured with the GRBM at 40–700 keV is 600 counts/s. The WFCdetermined position is α = 6^{h}51^{m}07^{s}, δ = +11^{o}35′.6 (equinox 2000.0), with an error radius of 6′ (99-percent confidence level). A follow-up observation with the Narrow Field Instruments is being carried out.”

COMET C/1999 H3 (LINEAR) Visual m_{1} estimates: Sept. 12.17 UT, 14.3 (C. S. Morris, Lockwood

Valley, CA, 0.51-m reflector); Oct. 12.75, 13.5 (M. Lehký, Hradec Králové, Czech Republic, 0.20-m reflector).

1999 October 15  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green